OpenStudy (anonymous):

Which lists all the integer solutions of the equation |x| = 4? A. –4 and 4 B. 0 and 4 C. –4 only D. 4 only

OpenStudy (anonymous):

@*Shadow*

OpenStudy (anonymous):

This one is confusing

OpenStudy (lyrae):

Do you know what absolute value is?

OpenStudy (anonymous):

Not really

OpenStudy (anonymous):

Lyrae, you got dis?

OpenStudy (lyrae):

Yeah!

OpenStudy (lyrae):

Well, it's actually quite simple. It means that the value between the || are turned non negative. So: |4| = 4 |-4| = 4 |100| = 100 |-100| = 100 Or generally |x| = x |-x| = x

OpenStudy (anonymous):

oh. So whatever is in the box, if its negative, outside the box its positive?

OpenStudy (lyrae):

Yeah, exactly!

OpenStudy (anonymous):

Cool, thanks, so would it be D?

OpenStudy (anonymous):

@Lyrae

OpenStudy (lyrae):

Nah, since the x is in the || and x can be any value positive or negative.

OpenStudy (anonymous):

so it would be A

OpenStudy (lyrae):

Yeah!

OpenStudy (anonymous):

Thanks alot

OpenStudy (anonymous):

I'll tag you if i have anymore problems

OpenStudy (lyrae):

Yw! :) Sure! I don't know how long I'll be online though... It's qute late here.

OpenStudy (anonymous):

What time is it? And where are you

OpenStudy (anonymous):

Australia and 2 AM?

OpenStudy (anonymous):

Europe and 11pm?

OpenStudy (lyrae):

Europe

OpenStudy (anonymous):

Oh wow

OpenStudy (anonymous):

11-2/3pm Time Frame?

OpenStudy (lyrae):

?

OpenStudy (lyrae):

But yeah, to get back on subject. When you solve equations like |x| = 4. All you usually have to do is to replace the || by a () and a -(). So |x| = 4 solution x = 4 -x = 4 => x = -4 Or if you have something slightly more complicated like |x-4| = 10 the soultion would be (x-4) = 10 => x = 14 -(x-4) = 10 => x = -6